OpenStudy (anonymous):

solve by substitution 2x + y = 9 y = 2x - 3

OpenStudy (anonymous):

Y = 9-2x 9-2x = 2x -3 4x = 12 x = 3 y = 9-6 y = 3

OpenStudy (anonymous):

Thank you